# Linear preservers of rank $k$ projections

### v1

@@ 1. Introduction
Let $\mathcal{H}$ be a complex Hilbert space. We say that a bounded linear operator on $\mathcal{H}$ is a projection if it is idempotent and self-adjoint. In this paper, we will study linear maps which send projections of a fixed finite rank to finite rank projections. It is well-known that if $k\in\mathbb{N}$, $k<\dim\mathcal{H}$, then the real linear span of the set of all rank $k$ projections $\mathcal{P}_{k}\left(\mathcal{H}\right)$ equals the real linear space of finite rank bounded self-adjoint operators $\mathcal{F}_{s}\left(\mathcal{H}\right)$, see e.g. [Molnar, Lemma 2.1.5.]. Thus, it makes sense to assume that the considered linear maps map from $\mathcal{F}_{s}\left(\mathcal{H}\right)$ to $\mathcal{F}_{s}\left(\mathcal{K}\right)$.
Our motivation to study such maps has roots in the famous Wigner’s theorem which plays a central role in mathematical foundations of quantum mechanics. It was published in 1931 and one can find the English translation in [wig]. The theorem can be formulated as follows.
@@
This theorem has been generalized in several ways. First, note that the assumption of injectivity is redundant since it clearly follows from (1.1). It turns out that neither surjectivity is needed in order to get an analogous result, only the isometry $U$ needs not be surjective, see e.g. [geh-wig].
Another way to generalize Wigner’s theorem is to describe preservers of a quantity on $\mathcal{P}_{k}\left(\mathcal{H}\right)$ which is equivalent to the transition probability $\operatorname{tr}PQ$ in the case $k=1$. Such an example is the description of surjective isometries on $\mathcal{P}_{k}\left(\mathcal{H}\right)$ with respect to the gap metric, i.e. the metric induced by the operator norm on $\mathcal{P}_{k}\left(\mathcal{H}\right)$, due to Gehér and Šemrl [gehse]. In Section 2, we will see that this metric is relevant in our study. Moreover, in the case $k=1$ it equals $\sin\varphi$, where $\varphi$ is the angle between the one-dimensional subspaces $\operatorname{ran}P$ and $\operatorname{ran}Q$, while $\operatorname{tr}PQ$ equals $\cos^{2}\varphi$. Here, $\operatorname{ran}A$ denotes the range of an operator $A$.
Let us introduce the result which is even more important for our study. The concept of an angle between one-dimensional subspaces can be generalized by the concept of principal angles between finite-dimensional subspaces. This was introduced by Jordan [jor] and is related to the Halmos’ two projections theorem. We will describe this topic more in Section 2. At this point, we mention Molnár’s characterization of preservers of principal angles.
Theorem 1.2 (L. Molnár, [mol1], [mol2]).
@@
    an example of injective $\tau\colon\mathcal{H}_{0,2k}\to\mathcal{U}_{m}$ satisfying (1.3) for $m=4k^{2}$. If $k\geq 2$ and $\phi$ is not an integer multiply of $\tfrac{\pi}{2}$, then $\tau$ does not preserve singular values, so it is not of the standard form from Proposition 3.4.
@@
Denote by $\mathcal{H}\otimes\mathcal{K}$ the tensor product of Hilbert spaces. If $A$ and $B$ are bounded linear operators on $A$ and $B$, respectively, then $A\otimes B$ denotes their tensor product. In the finite-dimensional case, $A\otimes B$ will denote the Kronecker product of matrices $A$ and $B$.
@@
  • If $\dim\mathcal{H}=n\in\mathbb{N}$ and $P_{0},Q_{0}\in\mathcal{P}_{f}\left(\mathcal{H}\right)$, then $\psi$ maps into $\mathcal{F}_{s}\left(\mathcal{H}\otimes\mathcal{K}\right)$ and $\psi\left(\mathcal{P}_{k}\left(\mathcal{H}\right)\right)\subset\mathcal{P}_{m}\left(\mathcal{H}\otimes\mathcal{K}\right)$, where $m=k\operatorname{rank}P_{0}+(n-k)\operatorname{rank}Q_{0}$.
@@
In [pp], Pankov and Plevnik considered what happens if one omits the injectivity assumption in Theorem 1.4. They showed that in the very special case $k=1$, the only additional resulting map is the one from Example 4. The main result of this paper is that the same conclusion holds in the general case, see Theorem 5.1. In order to prove this result, we will first prove Theorem 3.13 which is interesting in itself and characterizes linear maps $\mathcal{H}_{0,2k}\to\mathcal{H}_{0,2k}$ which preserve the set $\mathcal{I}\mathcal{H}_{0,2k}$. We will show that all such maps are unitary similarities, possibly composed by the entry-wise complex conjugation or multiplication by 1-1.
@@ 2. The set $\chi_{a}\left(X,Y\right)$
In [geh-grass, Lemma 2], it was characterized, when $\chi_{1}\left(X,Y\right)$ is a one-dimensional manifold for $X,Y\in\mathcal{G}_{k}\left(\mathcal{H}\right)$. The aim of this section is to precisely describe the set $\chi_{a}\left(X,Y\right)$ for $X,Y\in\mathcal{G}_{f}\left(\mathcal{H}\right)$. In order to understand this set better, we will use the Wedin’s version of the Halmos’ two projections theorem. One can obtain this formulation by applying [bsp, Corollary 2.2].
@@
Note that $X$ and $Y$ are subspaces of the finite-dimensional complex vector space $X+Y$. In this context, $\varphi_{1},\ldots,\varphi_{d}$ from Proposition 2.1 are precisely principal angles between $X$ and $Y$ from the interval $\left(0,\tfrac{\pi}{2}\right)$, while $\tfrac{\pi}{2}$ and 0 are principal angles of respective multiplicities $p+q$ and $m$. An interested reader can find more about principal angles e.g. in [Galantai], [bsp] or [bhat, Section VII.1].
@@
We will need the following property of solutions of the Sylvester equation. Denote the sets of bounded linear operators on $\mathcal{H}$ by $\mathcal{B}\left(\mathcal{H}\right)$. For any $A\in\mathcal{B}\left(\mathcal{H}\right)$, we denote its spectrum by $\sigma(A)$. Let $X\oplus Y$ be the direct sum of $X,Y\in\mathcal{G}\left(\mathcal{H}\right)$. Depending on the context, $A\oplus B$ may denote the direct sum of matrices or operators $A$ and $B$.
@@
for some $Q\in\mathcal{P}_{p+q}$, and $R_{j}\in\mathcal{P}_{m_{j}}$, $j=1,\ldots,r$, such that $0_{p+q}=(a-1)I_{p+q}$ and
@@ 3. Real linear maps transforming trace zero hermitian unitary matrices
The crucial results in this section are Proposition 3.11 and Theorem 3.13. In order to prove them, we need some technical lemmas
@@
Let $Y\in\mathcal{M}_{m}$ satisfy (3.2). Then $X^{\ast}X+Y^{\ast}Y=I_{m}$ and $\operatorname{Re}\left(X^{\ast}Y\right)=0_{m}$. Hence, $\|X\|\leq 1$, so $X$ has the above described singular value decomposition. Replace the pair $(X,Y)$ by the pair $(U^{\ast}XV^{\ast},U^{\ast}YV^{\ast})$ and denote
@@
Then there exist $U,V\in\mathcal{U}_{m}$, $p,q,r\in\mathbb{N}\cup\{0\}$, $n_{1},\ldots,n_{r}\in\mathbb{N}$, and real linear maps $\psi\colon\mathcal{H}_{0,2(k-1)}\to\mathcal{M}_{p}$ and $\phi_{j}\colon\mathcal{H}_{0,2(k-1)}\to\mathcal{H}_{n_{j}}$ such that $q+r>0$, $p+2q+2\sum_{j=1}^{r}n_{j}=m$, $\psi\left(\mathcal{I}\mathcal{H}_{0,2(k-1)}\right)\subset\mathcal{U}_{p}$, $\phi_{j}\left(\mathcal{I}\mathcal{H}_{0,2(k-1)}\right)\subset\mathcal{I}\mathcal{H}_{n_{j}}$, $j=1,\ldots,r$, and for all $A\in\mathcal{H}_{0,2(k-1)}$, $B\in\mathcal{H}_{0,2}$ we have
$\phi\left(A\oplus B\right)=U\left(\psi(A)\oplus\left(B\otimes I_{q}\right)\oplus\bigoplus_{j=1}^{r}\left(s_{j}B\otimes I_{n_{j}}+i\sqrt{1-s_{j}^{2}}I_{2}\otimes\phi_{j}(A)\right)\right)V.$
@@
$\phi\left(A\oplus 0_{2}\right)=Z\oplus 0_{2q}\oplus\left(\bigoplus_{j=1}^{r}\sqrt{1-s_{j}^{2}}Z_{j}\right),$ $A\in\mathcal{H}_{0,2(k-1)},$
@@
Assume now that $k\geq 2$ and that the statement holds for k1k-1. Let $U,V\in\mathcal{U}_{2k}$, $p,q,r$ and $n_{j}$, $j=1,\ldots,r$, be as in Lemma 3.8. After composing $\phi$ with a unitary similarity, we may assume that $V=I_{2k}$.
@@
Denote $\tau(T)=\begin{bmatrix}A^{\prime}&T^{\prime}\\ T^{\prime\ast}&B^{\prime}\end{bmatrix}$ for some $A^{\prime}\in\mathcal{H}_{2(k-1)}$, $B^{\prime}\in\mathcal{H}_{2}$, $T^{\prime}\in\mathcal{M}_{2(k-1),2}$. It follows from (3.8) that $K\circ B^{\prime}=0_{2}$, $K\in\mathcal{I}\mathcal{H}_{0,2}$. By Lemma 3.3, $B^{\prime}=0_{2}$. We now deduce from (3.7) that $A^{\prime 2}+T^{\prime}T^{\prime\ast}=TT^{\ast}$ and $T^{\prime\ast}T^{\prime}=I_{2}$. Hence, $\operatorname{tr}A^{\prime 2}=0$, implying that $A^{\prime}=0_{2(k-1)}$. We obtain from (3.8) that $sTKT^{\ast}T^{\prime}=T^{\prime}K$, $K\in\mathcal{I}\mathcal{H}_{0,2}$. Consequently, $sKT^{\ast}T^{\prime}=T^{\ast}T^{\prime}K$, $K\in\mathcal{I}\mathcal{H}_{0,2}$. By Lemma 3.3, $s=1$ and $T^{\ast}T^{\prime}=zI_{2}$ for some $z\in\mathbb{C}$. This and $T^{\prime}T^{\prime\ast}=TT^{\ast}$ imply that $T^{\prime}=zT$. We have shown that
@@ 4. Linear maps sending projections of a fixed finite rank to projections of a finite rank
for some $Q\in\mathcal{P}_{2p}$ and $U_{j}\in\mathcal{U}_{m_{j}}$, $j=1,\ldots,r$. Thus, there exists a unital linear map $\phi_{0}\colon\mathcal{H}_{2k}\to\mathcal{H}_{2p}$ such that $\phi_{0}(\mathcal{P}_{2k,k})\subset\mathcal{P}_{2p,p}$, and for every $j\in\left\{1,\ldots,r\right\}$ there exists a real linear map $\psi_{j}\colon\mathcal{H}_{2k}\to\mathcal{M}_{m_{j}}$ such that $\psi_{j}(\mathcal{P}_{2k,k})\subset\mathcal{U}_{m_{j}}$ and for each $A\in\mathcal{H}_{2k}$ we have
@@ 5. Linear preservers of rank $k$ projections
If $\dim\mathcal{H}=2k$, then the proof is complete. Assume now that $\dim\mathcal{H}>2k$. For $P,Q\in\mathcal{P}_{k}\left(\mathcal{H}\right)$ we denote $P\perp Q$ if $PQ=0$ (which is equivalent to $\operatorname{ran}P\perp\operatorname{ran}Q$). We have in particular shown that:
@@
It remains to show that $\phi(P)=P_{0}$, $P\in\mathcal{P}_{k}\left(\mathcal{H}\right)$. So, let $P\in\mathcal{P}_{k}\left(\mathcal{H}\right)\setminus\left\langle W_{0}\right]$. Then there exist $Q,R\in\mathcal{P}_{k}\left(\mathcal{H}\right)\cap\left\langle W_{0}\right]$ such that $P\perp R$ and $Q\perp R$. Let $P^{\prime}\in\mathcal{P}_{k}\left(\mathcal{H}\right)$ such that $\operatorname{ran}P^{\prime}\subset\operatorname{ran}P+\operatorname{ran}Q$, $P^{\prime}\neq Q$, and $P^{\prime}\not\perp Q$. The latter yields $\phi(P^{\prime})\not\perp\phi(Q)=\phi(R)$. Since $P^{\prime}\perp R$, this implies $\phi(P^{\prime})=\phi(R)=P_{0}$. If $\operatorname{ran}P\not\perp W_{0}$, then we can choose $Q$ and $R$ above such that $P\not\perp Q$, so we can choose $P^{\prime}=P$. If $\operatorname{ran}P\perp W_{0}$, then $P\perp Q$ and $P+Q-P^{\prime}$ fulfills the above assumptions of $P^{\prime}$. Consequently, $\phi(P)=\phi(P+Q-P^{\prime})=P_{0}$.
@@
see e.g. [faure, Theorem 4.1], yielding that $\phi(A)=VAV^{\ast}$, $A\in\mathcal{F}_{s}\left(\mathcal{H}\right)$. ∎

### v2

@@ 1. Introduction
Let $\mathcal{H}$ be a complex Hilbert space. We say that a bounded linear operator on $\mathcal{H}$ is a projection if it is idempotent and self-adjoint. In this paper, we will study linear maps which send projections of a fixed finite rank to finite rank projections. It is well-known that if $k\in\mathbb{N}$, $k<\dim\mathcal{H}$, then the real linear span of the set of all rank $k$ projections $\mathcal{P}_{k}\left(\mathcal{H}\right)$ equals the real linear space of finite rank bounded self-adjoint operators $\mathcal{F}_{s}\left(\mathcal{H}\right)$, see e.g. [9, Lemma 2.1.5.]. Thus, it makes sense to assume that the considered linear maps map from $\mathcal{F}_{s}\left(\mathcal{H}\right)$ to $\mathcal{F}_{s}\left(\mathcal{K}\right)$.
Our motivation to study such maps has roots in the famous Wigner’s theorem which plays a central role in mathematical foundations of quantum mechanics. It was published in 1931 and one can find the English translation in [14]. The theorem can be formulated as follows.
@@
This theorem has been generalized in several ways. First, note that the assumption of injectivity is redundant since it clearly follows from (1.1). It turns out that neither surjectivity is needed in order to get an analogous result, only the isometry $U$ needs not be surjective, see e.g. [5].
Another way to generalize Wigner’s theorem is to describe preservers of a quantity on $\mathcal{P}_{k}\left(\mathcal{H}\right)$ which is equivalent to the transition probability $\operatorname{tr}PQ$ in the case $k=1$. Such an example is the description of surjective isometries on $\mathcal{P}_{k}\left(\mathcal{H}\right)$ with respect to the gap metric, i.e. the metric induced by the operator norm on $\mathcal{P}_{k}\left(\mathcal{H}\right)$, due to Gehér and Šemrl [6]. In Section 2, we will see that this metric is relevant in our study. Moreover, in the case $k=1$ it equals $\sin\varphi$, where $\varphi$ is the angle between the one-dimensional subspaces $\operatorname{ran}P$ and $\operatorname{ran}Q$, while $\operatorname{tr}PQ$ equals $\cos^{2}\varphi$. Here, $\operatorname{ran}A$ denotes the range of an operator $A$.
Let us introduce the result which is even more important for our study. The concept of an angle between one-dimensional subspaces can be generalized by the concept of principal angles between finite-dimensional subspaces. This was introduced by Jordan [8] and is related to the Halmos’ two projections theorem. We will describe this topic more in Section 2. At this point, we mention Molnár’s characterization of preservers of principal angles.
Theorem 1.2 (L. Molnár, [10], [11]).
@@
    is an example of injective $\tau\colon\mathcal{H}_{0,2k}\to\mathcal{U}_{m}$ satisfying (1.3) for $m=4k^{2}$. If $k\geq 2$ and $\varphi$ is not an integer multiple of $\tfrac{\pi}{2}$, then $\tau$ does not preserve singular values, so it is not of the standard form from Proposition 3.4.
@@
Denote by $\mathcal{H}\otimes\mathcal{K}$ the tensor product of Hilbert spaces. If $A$ and $B$ are bounded linear operators on $\mathcal{H}$ and $\mathcal{K}$, respectively, then $A\otimes B$ denotes their tensor product. In the finite-dimensional case, $A\otimes B$ will denote the Kronecker product of matrices $A$ and $B$.
@@
  • If $\dim\mathcal{H}=n\in\mathbb{N}$ and $P_{0},Q_{0}\in\mathcal{P}_{f}\left(\mathcal{K}\right)$, then $\psi$ maps into $\mathcal{F}_{s}\left(\mathcal{H}\otimes\mathcal{K}\right)$ and $\psi\left(\mathcal{P}_{k}\left(\mathcal{H}\right)\right)\subset\mathcal{P}_{m}\left(\mathcal{H}\otimes\mathcal{K}\right)$, where $m=k\operatorname{rank}P_{0}+(n-k)\operatorname{rank}Q_{0}$.
@@
In [12], Pankov and Plevnik considered what happens if one omits the injectivity assumption in Theorem 1.4. They showed that in the very special case $k=1$, the only additional resulting map is the one from Example 4. The main result of this paper is that the same conclusion holds in the general case, see Theorem 5.1. In order to prove this result, we will first prove Theorem 3.13 which is interesting in itself and characterizes linear maps $\mathcal{H}_{0,2k}\to\mathcal{H}_{0,2k}$ which preserve the set $\mathcal{I}\mathcal{H}_{0,2k}$. We will show that all such maps are unitary similarities, possibly composed by the entry-wise complex conjugation or multiplication by 1-1.
@@ 2. The set $\chi_{a}\left(X,Y\right)$
In [7, Lemma 2], it was characterized, when $\chi_{1}\left(X,Y\right)$ is a one-dimensional manifold for $X,Y\in\mathcal{G}_{k}\left(\mathcal{H}\right)$. The aim of this section is to precisely describe the set $\chi_{a}\left(X,Y\right)$ for $X,Y\in\mathcal{G}_{f}\left(\mathcal{H}\right)$. In order to understand this set better, we will use the Wedin’s version of the Halmos’ two projections theorem. One can obtain this formulation by applying [2, Corollary 2.2].
@@
Note that $X$ and $Y$ are subspaces of the finite-dimensional complex vector space $X+Y$. In this context, $\varphi_{1},\ldots,\varphi_{d}$ from Proposition 2.1 are precisely principal angles between $X$ and $Y$ from the interval $\left(0,\tfrac{\pi}{2}\right)$, while $\tfrac{\pi}{2}$ and 0 are principal angles of respective multiplicities $p+q$ and $m$. An interested reader can find more about principal angles e.g. in [4], [2] or [1, Section VII.1].
@@
We will need the following property of solutions of the Sylvester equation. Denote the set of bounded linear operators on $\mathcal{H}$ by $\mathcal{B}\left(\mathcal{H}\right)$. For any $A\in\mathcal{B}\left(\mathcal{H}\right)$, we denote its spectrum by $\sigma(A)$. Let $X\oplus Y$ be the direct sum of $X,Y\in\mathcal{G}\left(\mathcal{H}\right)$. Depending on the context, $A\oplus B$ may denote the direct sum of matrices or operators $A$ and $B$.
@@
for some $Q\in\mathcal{P}_{p+q}$, and $R_{j}\in\mathcal{P}_{2m_{j}}$, $j=1,\ldots,r$, such that $0_{p+q}=(a-1)I_{p+q}$ and
@@ 3. Real linear maps transforming trace zero hermitian unitary matrices
The crucial results in this section are Proposition 3.11 and Theorem 3.13. In order to prove them, we need some technical lemmas.
@@
Let $Y\in\mathcal{M}_{m}$ satisfy (3.2). Then $X^{\ast}X+Y^{\ast}Y=I_{m}$ and $\operatorname{Re}\left(X^{\ast}Y\right)=0_{m}$. Hence, $\|X\|\leq 1$, so $X$ has the above described singular value decomposition. Replace the pair $(X,Y)$ by the pair $(U^{\ast}XV^{\ast},U^{\ast}YV^{\ast})$ and define
@@
Then there exist $U,V\in\mathcal{U}_{m}$, $W\in\mathcal{U}_{2k}$, $p,q,r\in\mathbb{N}\cup\{0\}$, $n_{1},\ldots,n_{r}\in\mathbb{N}$, and real linear maps $\psi\colon\mathcal{H}_{0,2(k-1)}\to\mathcal{M}_{p}$ and $\phi_{j}\colon\mathcal{H}_{0,2(k-1)}\to\mathcal{H}_{n_{j}}$ such that $q+r>0$, $p+2q+2\sum_{j=1}^{r}n_{j}=m$, $\psi\left(\mathcal{I}\mathcal{H}_{0,2(k-1)}\right)\subset\mathcal{U}_{p}$, $\phi_{j}\left(\mathcal{I}\mathcal{H}_{0,2(k-1)}\right)\subset\mathcal{I}\mathcal{H}_{n_{j}}$, $j=1,\ldots,r$, and for all $A\in\mathcal{H}_{0,2(k-1)}$, $B\in\mathcal{H}_{0,2}$ we have
$\phi\left(W^{\ast}\left(A\oplus B\right)W\right)=U\left(\psi(A)\oplus\left(B\otimes I_{q}\right)\oplus\bigoplus_{j=1}^{r}\left(s_{j}B\otimes I_{n_{j}}+i\sqrt{1-s_{j}^{2}}I_{2}\otimes\phi_{j}(A)\right)\right)V.$
@@
$\phi\left(H\oplus 0_{2}\right)=Z\oplus 0_{2q}\oplus\left(\bigoplus_{j=1}^{r}\sqrt{1-s_{j}^{2}}Z_{j}\right)$
@@
Assume now that $k\geq 2$ and that the statement holds for k1k-1. Let $U,V,W\in\mathcal{U}_{2k}$, $p,q,r$ and $n_{j}$, $j=1,\ldots,r$, be as in Lemma 3.8. After replacing $\phi$ with the map $A\mapsto V\phi\left(W^{\ast}AW\right)V^{\ast}$, we may assume that $V=W=I_{2k}$.
@@
Denote $\tau(T)=\begin{bmatrix}A^{\prime}&T^{\prime}\\ T^{\prime\ast}&B^{\prime}\end{bmatrix}$ for some $A^{\prime}\in\mathcal{H}_{2(k-1)}$, $B^{\prime}\in\mathcal{H}_{2}$, $T^{\prime}\in\mathcal{M}_{2(k-1),2}$. It follows from (3.8) that $K\circ B^{\prime}=0_{2}$, $K\in\mathcal{I}\mathcal{H}_{0,2}$. By Lemma 3.3, $B^{\prime}=0_{2}$. We now deduce from (3.7) that $A^{\prime 2}+T^{\prime}T^{\prime\ast}=TT^{\ast}$ and $T^{\prime\ast}T^{\prime}=I_{2}$. Hence, $\operatorname{tr}A^{\prime 2}=0$, implying that $A^{\prime}=0_{2(k-1)}$. We obtain from (3.8) that $sTKT^{\ast}T^{\prime}=T^{\prime}K$, $K\in\mathcal{I}\mathcal{H}_{0,2}$. Consequently, $sKT^{\ast}T^{\prime}=T^{\ast}T^{\prime}K$, $K\in\mathcal{I}\mathcal{H}_{0,2}$. By Lemma 3.3, $s=1$ and $T^{\ast}T^{\prime}=zI_{2}$ for some $z\in\mathbb{C}$. This and $T^{\prime}T^{\prime\ast}=TT^{\ast}$ imply that $|z|=1$ and $T^{\prime}=zT$. We have shown that
@@ 4. Linear maps sending projections of a fixed finite rank to projections of a finite rank
for some $Q\in\mathcal{P}_{2p}$ and $U_{j}\in\mathcal{U}_{m_{j}}$, $j=1,\ldots,r$. By Proposition 4.1, $Q\in\mathcal{P}_{2p,p}$. Thus, there exists a unital linear map $\phi_{0}\colon\mathcal{H}_{2k}\to\mathcal{H}_{2p}$ such that $\phi_{0}(\mathcal{P}_{2k,k})\subset\mathcal{P}_{2p,p}$, and for every $j\in\left\{1,\ldots,r\right\}$ there exists a real linear map $\psi_{j}\colon\mathcal{H}_{2k}\to\mathcal{M}_{m_{j}}$ such that $\psi_{j}(\mathcal{P}_{2k,k})\subset\mathcal{U}_{m_{j}}$ and for each $A\in\mathcal{H}_{2k}$ we have
@@ 5. Linear preservers of rank $k$ projections
If $\dim\mathcal{H}=2k$, then the proof is complete. Assume now that $\dim\mathcal{H}>2k$. For $P,Q\in\mathcal{P}_{k}\left(\mathcal{H}\right)$ we write $P\perp Q$ if $PQ=0$ (which is equivalent to $\operatorname{ran}P\perp\operatorname{ran}Q$). We have in particular shown that:
@@
It remains to show that $\phi(P)=P_{0}$, $P\in\mathcal{P}_{k}\left(\mathcal{H}\right)$. So, let $P\in\mathcal{P}_{k}\left(\mathcal{H}\right)\setminus\left\langle W_{0}\right]$. Then there exist $Q,R\in\mathcal{P}_{k}\left(\mathcal{H}\right)\cap\left\langle W_{0}\right]$ such that $P\perp R$ and $Q\perp R$. Let $P^{\prime}\in\mathcal{P}_{k}\left(\mathcal{H}\right)$ such that $\operatorname{ran}P^{\prime}\subset\operatorname{ran}P+\operatorname{ran}Q$, $P^{\prime}\neq Q$, and $P^{\prime}\not\perp Q$. The latter yields $\phi(P^{\prime})\not\perp\phi(Q)=\phi(R)$. Since $P^{\prime}\perp R$, this implies $\phi(P^{\prime})=\phi(R)=P_{0}$. If $\operatorname{ran}P\not\perp W_{0}$, then we can choose $Q$ and $R$ above such that $P\not\perp Q$, so we can set $P^{\prime}$ to be $P$. If $\operatorname{ran}P\perp W_{0}$, then $P\perp Q$ and $P+Q-P^{\prime}$ fulfills the above assumptions on $P^{\prime}$. Consequently, $\phi(P)=\phi(P+Q-P^{\prime})=P_{0}$.
@@
see e.g. [3, Theorem 4.1], yielding that $\phi(A)=VAV^{\ast}$, $A\in\mathcal{F}_{s}\left(\mathcal{H}\right)$. ∎